### v1

# Ultra-stretchable and Self-Healable Vitrimers with Tuneable Damping and Mechanical Response

Abstract
Vitrimers are a relatively new class of polymer materials with unique properties offered by cross-links that can undergo associative exchange dynamics. We here present a new class of vitrimers based on poly(methyl acrylate) with cross-links utilising dioxaboralane metathesis. These vitrimers demonstrate a combination of ultra-stretchability (up to 80 times their own length), mechanical toughness ( 40 MJ/m3), and thermal stability up to 250 °C; moreover, the vitrimers demonstrate excellent mechanical damping characterised by a loss factor () with a maximum of 2–3 and an effective value 0.3 across five decades in frequency (0.001–100 Hz), or correspondingly across a -range of 35 °C near room temperature (for a probe frequency of 1 Hz). The vitrimers can be successfully re-processed using both a thermo-mechanical and a chemical processing route, and can self-heal at room temperature, making them suitable for sustainable applications. The material properties are directly tuneable by variation of both the amount of cross-linker and by the degree of curing. Thus, this class of vitrimers are promising for applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required.
I Introduction
Elastomers that combine considerable stretchability with high mechanical strength and toughness are important for a wide range of applications including actuators, sensors, soft robotics, wearable electronics, and materials for energy storage. [1, 2, 3] Standard elastomers, such as vulcanised rubbers, can readily achieve tensile strains of several 100% combined with tensile strengths of 10 MPa. [4, 5, 6] However, elastomers that can sustain strains of several 1000% combined with a high tensile strength (a high mechanical toughness) are challenging to produce. A variety of approaches have been attempted often focused on combining an elastic stress-bearing network with a network of sacrificial bonds that can dissipate the applied mechanical energy by bond fracture. Reported polymer systems include: block-copolymer-based thermoplastic elastomers, [7] interpenetrating double polymer networks, [8, 9] elastomers based on supramolecular interactions, [10, 11] or polymeric networks based on dynamic covalent crosslinks, so called covalent adaptable networks (CANs) [12, 13, 14, 15, 16]. However, elastomers that show very high stretchability ( 1000–10000%), typically also exhibit significant strain-softening upon yielding, leading to a rapidly decreasing mechanical strength upon deformation. [10, 12, 14, 17, 18, 19, 20, 21] Thus, it remains an on-going challenge to produce elastomers that combine high stretchability, mechanical strength and toughness.
Here, we demonstrate that a relatively new type of dynamically cross-linked networks, so called vitrimers, can achieve these requirements, [22, 23, 24], while also providing other attractive properties such as effective self-healing, excellent mechanical damping and both thermo-mechanical and chemical re-processability.
Polymers with permanent crosslinks – thermosets – show good mechanical and thermal stability, but poor flexibility, malleability and reprocessing capability. Conversely, polymers that lack permanent crosslinks – thermoplastics – offer reprocessability but lower stability. CANs [25] combine the advantageous properties of thermosets and thermoplastics, resulting in both mechanical stability and effective materials reprocessing. [26] The subset of CANs for which the crosslink bond exchanges are associative, are termed vitrimers. [22, 23, 24] Vitrimers keep their crosslink density fixed during bond exchange, which results in excellent mechanical and thermal stability. At high temperatures, these materials are typically malleable and thus reprocessable, [26] whereas at low temperatures, they show behaviour akin to permanently crosslinked thermosets with good mechanical and thermal stability. Vitrimers have been produced based on different strategies for the covalent exchange, such as: transesterification, [22] transcarbonation, [27] olefin metathesis, [28], transimination, [29] silyl ether metathesis, [30] or dioxaborolane metathesis. [31]
The unique properties of vitrimers suggest that they should be ideal materials to combine properties such as stretchability and material strength with efficient re-processing ability. [22, 31, 14] Moreover, the dynamic nature of the bond exchange should be important in achieving other important properties such as good self-healing [32], material welding [33] capabilities, and high mechanical dissipation. [34, 35] Thus, vitrimers are promising and exciting candidates for the next generation of polymer-based applications, [36] including synthetic polymers acting as artificial skin, muscles, or cartilage that combine high stretching and self-healing ability, with large mechanical strength and high mechanical dissipation. [1, 2, 3, 37, 38, 36]
Here, we present a new ultra-stretchable vitrimer system based on poly(methyl acrylate) (PMA), where the PMA-based vitrimers are prepared by free radical polymerization (FRP) of the methyl acrylate monomer and a diallyl dioxaborolane dynamic crosslinker. The associative crosslink exhange reaction, based on dioxaboralane metathesis, shows excellent thermal stability and a relatively low activation energy (16 kJ/mol), which allows rapid bond exchange without added solvent, catalyst or free diols. [31] Long chain-length PMAs are viscous fluids at room temperature, characterised by glass-transition temperatures 15°C, and are often used in coatings and adhesives, as liquid crystal elastomer backbone components, or as components in block co-polymers to improve material flexibility. [39, 40, 41] Here, PMA was chosen as the matrix polymer both due to its lack of crystallinity and due to its flexible nature, as reflected in a relatively low glass transition temperature, which ensures a high segmental mobility.
Our vitrimers demonstrate stretchability of up to 8000, combined with a high tensile strength ( 0.6 MPa) and toughness ( 38.5 MJ/m3). They also demonstrate room-temperature self-healing and a high mechanical loss factor with tan() 1 over three orders of magnitude, and tan() 0.3 over five orders of magnitude, in frequency (– Hz), as well as excellent re-processing through either a thermomechanical (hot-press) or a chemical route based on hydrolysis of the boronic ester. Thus, our PMA-based vitrimers are promising materials for a wide range of application areas.
II Results and Discussion
II.1 Vitrimer synthesis and characterization
A set of poly(methyl acrylate) based vitrimers (VMAs) of varying crosslink densities were produced, as shown schematically in Figure 1(a); the dynamic crosslinker was first synthesized, [42] and subsequently copolymerized with the methyl acrylate monomer using FRP to produce the vitrimers. The schematic in Figure 1(b) illustrates how the associative bond exchange results in topological rearrangements, which in turn lead to malleability, even though the crosslink density is maintained. After synthesis, the samples were dried in an oven and then hot-pressed at °C, to produce vitrimers of nominal crosslink density and mol% (by stochiometry). 1H NMR spectra of the reactants (monomer and crosslinker) and the synthesis product are shown in Figure 1(c). The successful polymerization of methyl acrylate is demonstrated by the disappearance of signals due to the monomer double bond (peaks a-c in Figure 1(c)). The NMR spectra demonstrate a 50-60% incorporation of the crosslinker after the initial FRP step (peaks labelled 1-3 in Figure 1(c)). The lower (than 100%) incorporation is due to the lower reactivity of the double bond in the allyl compared to the acrylate moiety. [43] We refer to these first-cured samples as semi-cured and denote them as VMA-’ (the apostrophe marks that they are semi-cured and denotes the nominal molar cross-link density). To produce fully cured vitrimers, samples were further heated to °C for 2 hours under vacuum. After this process, no peaks due to the allyl double bond are observed in NMR, confirming that all vinyl bonds were fully incorporated; we thus refer to these samples as fully cured and denote them as VMA-. 1H NMR studies on the fully cured vitrimers (shown in Figure 2(a)) demonstrate that the resonance at 7.72 ppm, assigned to the phenyl ring of p-phenylene diboronic acid after hydrolysis (see inset), increases systematically with the nominal crosslink density, as expected for full incorporation of the crosslinker.
To characterise the produced chain-length distribution of the vitrimer samples, a set of samples were produced for which the crosslinks were removed by hydrolysis and characterised by gel permeation chromatography (GPC) (see the Experimental Section for details). The GPC traces and results are shown in Figure S1 and Table S1, demonstrating that the pure polymers all have number-averaged molecular weights close to 20 kDa and broad dispersities 4. The entanglement molecular weight for pure linear PMA is 11kDa, and the critical molecular weight is 20 kDa[44]. Thus, based on our GPC results, we expect the vitrimers to be weakly entangled and characterised by a broad distribution of chain-lengths. Representative photos of the produced vitrimers are included in Figure 1(d), which shows the mol % semicured (VMA-1’) and fully-cured (VMA-1) samples, respectively, together with a photo illustrating that a strip of VMA-1’ (width 4.0 × thickness 0.8 mm) can readily lift a weight of 500 g, thus demonstrating the mechanically robust nature of the produced vitrimers.
Differential scanning calorimetry (DSC) was performed to further characterise the vitrimers. The temperature () dependent heat flow recorded for a heating rate of 10 °C/min is shown in Figure 2(b) for two representative vitrimer samples (VMA-1’ and VMA-1). Both the glass transition temperature () and the glass transition breadth () increase upon the additional curing (from VMA-1’ to VMA-1). DSC data for all vitrimer samples are shown in Figure S2 and the corresponding analysis results are shown in Table S2. Consistent with the results shown in Figure 2(b), both and increase upon additional curing for all cross-link densities. Generally, an increase in the crosslink density (due to an increase in the nominal crosslink density or due to additional curing) restricts the polymer mobility, thus resulting in increased values, as shown in Figure S2. Similar behaviour has previously been observed for thermosets, [45] supramolecular polymers, [46] and vitrimers. [29, 47] However, we note that for our semicured vitrimers, dangling or ‘free’ crosslinkers might act as plasticizers that can at least partially counteract the increase, as shown in Figure S2.
Thermogravimetric analysis (TGA) was performed to determine the vitrimer thermal stability. Data for all vitrimer samples, together with data for a linear PMA with a degree of polymerization of ( g/mol), are shown in Figure 2(c). The TGA data can be roughly divided into three regimes: Regime I: stable with no significant weight loss (250 °C); Regime II: undergoing thermal degradation, including main-chain scission, [48] as previously demonstrated for PMA ( °C); Regime III: degradation is practically complete ( °C). The temperature , denoting a 5% thermal weight loss, is increased by 30°C between the linear PMA and the PMA-based vitrimers.
II.2 Vitrimer Reprocessability
A strong advantage of vitrimers is the thermal reprocessability afforded by the dynamic nature of the cross-links [22, 31, 49]. Alternative chemical reprocessing routes, either by depolymerization or by degradation (e.g. through high temperature or acidic conditions), have also been demonstrated for vitrimers; [27, 50] however, these routes breaks down the vitrimer structure and reprocessing thus involves partially repeating the synthesis. In contrast, our vitrimers can be effectively re-processed using either a thermo-mechanical or a chemical route. For thermal reprocessing, the vitrimer samples were cut into pieces and placed into a hot-press mould, as shown in Figure 3(a)), where they were kept for 15 minutes at °C to facilitate dynamic cross-link exchange, and thus reprocessability. In contrast, the solvent-based reprocessing route is based on boronic ester hydrolysis, where the solvent needs to contain a sufficient amount of water for the hydrolysis (we used acetone mixed with a small amount of water; vol/vol = 200:1). The presence of the solvent initially leads to swelling of the vitrimers, followed by dissolution, where water acts to chemically “unzip” the cross-links (see the sketch in Figure 3(a)); NMR characterisation demonstrated that the vitrimer samples could be successfully hydrolysed using this approach, as shown in Figure 2(a). Following the dissolution process, the vitrimers were recovered by solvent removal in a vacuum oven at 80 °C, after which the dry samples were hot-pressed into samples of suitable dimensions. To investigate how successful each reprocessing route is, tensile stress-strain measurements were performed both before and after reprocessing. As shown in Figure 3(b) for VMA-1’, both the tensile strengths and elongation at break are largely recovered for samples undergoing either of the two reprocessing routes (2 MPa and 4500 %), and similar behaviour is observed for the corresponding fully cured sample (VMA-1), as shown in Figure S3.
II.3 Vitrimer Network Response and Relaxation
To evaluate the rheological response of the crosslinked network, we performed small-amplitude oscillatory shear (SAOS) rheology. The SAOS measurements were carried out from 15 °C to 240 °C, and master curves for the storage and loss shear moduli ( and ) were produced based on the time-temperature superposition (TTS) principle, as shown in Figure 4 and described in detail in the Experimental Section. TTS could be accurately performed both at low and high , where the shift parameters obtained for each -range refer to different relaxation behaviour. For the low data, we choose a reference temperature 20°C for which the structural relaxation response (crossover of and ) lies within the experimental window. For the high data, we instead choose a reference temperature of 240°C for which the onset of the high- relaxation behaviour lies within the experimental window.
For each of the four vitrimer samples (VMA-1 to VMA-8; see Figure 4(a-d)) a glassy response is observed at the highest frequencies corresponding to a peak in and the establishment of a glassy plateau in ; for lower frequencies, a spectrum of Rouse modes are observed, as indicated by , followed by a plateau in and a corresponding minimum in . We define the plateau moduli () as the value corresponding to the minima in , following a commonly used approach in the literature, [51], as shown by the arrows in Figure 4(a-d). As the nominal cross-link density increases, the plateau becomes flatter and thus more well-defined and correspondingly increases in an approximately linear manner, as shown in Figure 4(f) (green circles). We note that particularly for the lowest crosslink density sample (VMA-1), the ‘plateau’ behaviour is poorly defined demonstrating the presence of a broad spectrum of relaxation modes across 8 orders of magnitude in frequency; this is consistent with the behaviour of a weakly cross-linked (and weakly entangled) system, which contains a broad distribution of dangling polymer strands, or free polymer chains, thus resulting in a very broad spectrum of relaxation frequencies. Finally, for the lowest frequencies (240 °C) an additional relaxation regime is clearly observed, as shown by the crossover of the storage and loss moduli.
The plateau modulus can generally be attributed both to the covalent dynamic crosslinks and to topological crosslinks due to chain entanglements. In Figure S4, we compare the crosslink density , as determined from the plateau modulus (), with the nominal crosslink density determined from the chemical composition, showing that and demonstrating an approximately linear relation, even though a significant difference is observed between the magnitude of the nominal and actual crosslink density. The average molecular weight between crosslinks () can be estimated from , [52] where is the mass density 0.8 g/cm3 at 160 °C for VMA (see the detailed discussion in the Experimental Section). The estimated for VMA- (1, 2.5, 5 and 8) are: 80 kDa, 14 kDa, 6 kDa, and 4 kDa, respectively. For comparison, the entanglement molecular weight () for linear poly(methyl acrylate) (PMA) is 11 kDa, [44], so that is larger than at low crosslink density, and smaller than at high crosslink density. For context, it is also important to compare both and with the length of the highly polydisperse primary chains given in Table S1. For VMA-1 and VMA-2.5 we find ; for VMA-5 and VMA-8 we find . In all cases, due to the high polydispersity, we thus expect a significant number of chains that have only a small number of dynamic crosslinks, and/or which are shorter than the entanglement molecular weight.
Both with regards to the significant difference between and , and the small influence of entanglements in determining the ‘rubbery’ plateau, we note that: (i) since the reactivity of the crosslinker is much lower than that of the acrylate monomer, we expect a degree of heterogeneity in the cross-link distribution, and the formation of “effective cross-links” (each containing several nominal cross-links) is not surprising; (ii) the calculation of the nominal cross-link density does not take account of the possibility of cross-link formation within the same chain; such ‘loops’ are not generally stress-bearing and will thus not contribute to the plateau modulus; (iii) it is highly likely, especially at low crosslink density, that many of the crosslinks lie within branched structures that are either “free” chains wholly unconnected to the network, or “side branches” connected to the network at one point only: such structures can relax their internal configuration, and so relax their contribution to the stress, without breaking of crosslinks and thus do not contribute to the plateau; (iv) finally, relaxation of such free chains or side branches will release entanglements through the well known process of dynamic dilution [51], thus progressively increasing the effective entanglement molecular weight and reducing the influence of entanglements on the plateau.
To further characterise the rheological response, we define two characteristic relaxation times from the crossover frequencies between the storage and loss moduli. The high-frequency (low-) relaxation time is defined as , and the low frequeny (high-) relaxation time as , where near the former corresponds to the structural relaxation, and the latter corresponds to the relaxation due to dynamic cross-link exchange. denotes the characteristic angular frequency of the low- and of the high- relaxation (each determined at their corresponding reference temperatures), and and are the two horizontal shift parameters, respectively. for the four fully cured vitrimer samples are shown in an Arrhenius plot in Figure S5 demonstrating only very small differences with a variation of the cross-link density. The corresponding are shown in Figure S6 together with the values determined from DSC, showing a slowing down for increasing crosslink density consistent with the observed increase in .
The relaxation times are shown in an Arrhenius plot in Figure 4(e). For each of the four samples, demonstrate Arrhenius behaviour: ( denotes a particular sample), where is the gas constant, as shown by the fits (dashed lines). The activation enthalpy (), determined for each of the four fits, are plotted in Figure 4(f) (squares). We find that is of similar magnitude 80 kJ/mol for the three higher crosslink density samples, whereas a somewhat lower of kJ/mol is found for the 1 mol % sample. These values are slightly higher than those previously reported for vitrimers based on the same dynamic crosslinker, where examples include: polysiloxane vitrimers (20–57 kJ/mol), [53] vitrimers prepared by a thiol-ene “click” reaction (50 kJ/mol), [54] polyolefin vitrimers (8–47 kJ/mol), [55] and poly(styrene-b-isoprene-b-styrene) vitrimers (44 kJ/mol). [56] As shown in Figure 4(f) the activation energy increases monotonically with increasing nominal crosslink density, even though the increase for crosslink densities above 2.5 mol % is very small; this behaviour is consistent with reports for vitrimers based on both dioxaboralane and other exchange mechanisms. [29, 57]
Finally, a brief comparison with the behaviour of semi-cured vitrimers is included in Figure S7 (SI) that shows the TTS mastercurves for the semi-cured VMA-1’. At high frequencies, the relaxation response is observed, followed by a plateau-like behaviour that is less distinct than the one observed for the corresponding fully cured sample (VMA-1) and the low frequency (high-) relaxation is much less pronounced. The -dependent relaxation times are shown for both the VMA-1’ and the VMA-1 sample in Figure S8 (SI). The relaxation time of VMA-1’ is faster than that of VMA-1, consistent with the fact that of VMA-1’ is lower than that of VMA-1. Moreover, the relaxation times for the slow network relaxations are also shown. Relatively small differences are observed in the activation enthalpies, ranging from 69 kJ/mol for VMA-1’ to 56 kJ/mol for VMA-1, and the most significant observed change is the shift of towards shorter time-scales for the semi-cured sample, which is consistent with the increase in chain mobility for the less cross-linked VMA-1’ sample.
II.4 Tensile deformation and stretchability
The room temperature ( = 20 °C) tensile stress response () to an imposed strain () was determined for four different extension speeds of 10, 50, 100 and 200 mm/min. Data for the semi-cured VMA-1’ samples are shown in Figure 5(a) as engineering stress () versus imposed strain. For small deformations, a linear elastic regime is observed for all applied strain rates, corresponding to a Young’s modulus of 50 MPa, as shown in the inset to Figure 5(a). For larger deformations, the observed yield stress increases with applied strain rate within the range 0.4–1.1 MPa, and strain hardening is observed where the stress required for deformation rises more rapidly for increasing strain rates. Such complex strain-rate dependence is typically observed both in linear polymers [58] and in elastomers based on dynamically crosslinked networks. [14, 59, 60, 61]
For the lowest applied strain rate, only a very weak strain hardening effect is found and a nearly horizontal plateau is observed, whereas higher rates lead to an increasing growth of the stress response. For an elastomer that is capable of large tensile strains, the engineering strain, which assumes a constant cross-sectional area, is not always representative of the tensile strength. Thus, in the inset to Figure 5(a) we also present the estimated true stress () by assuming a constant volume during stretching. For the three lowest extension speeds, the true stress at break is effectively constant at 80 MPa, whereas at the slowest deformation, the true stress reaches only 50 MPa – hence, deformation at a faster rate results in the vitrimer reaching its ‘true’ strength at break, and thus fracture, sooner.
Macroscopic failure of materials commonly starts with micro-cracks, where crack propagation leads to the final failure. [69] For vitrimers, both a sufficient segmental mobility (low enough ) and any associative bond kinetics fast enough to allow network relaxations on the relevant deformation time-scale [61] could deter these mechanisms and thus contribute to large stretchability. Since °C is close to room temperature, the characteristic time-scale of the structural () relaxation is relatively slow, and 0.1 s (see Fig. S8). As an example for the four semi-cured samples, the imposed extensional strain rates () are 0.0267, 1.691, 2.251, and 3.024 s-1, respectively. In turn, the time-scales corresponding to the inverse applied strain rates () are 37.5, 0.59, 0.49, and 0.33 s, demonstrating that the time-scale characteristic of the deformation is for the three higher strain rates, whereas for the lowest applied rate. Thus, particularly for the slowest deformation rate, there is enough time for structural relaxation to take place during the deformation. It is also possible that the rate of the relaxation and/or the bond exchange mechanism is significantly sped up, particularly where the stress is focused near the tip of any micro-cracks, resulting in further contributions to these effects.
The overall result is a negligible strain-hardening for the lowest applied strain rate and a remarkably stretchable material, for which a deformation at break of 8000% is observed; the corresponding deformation at break is reduced to 2500% for the faster rate of 3.024 s-1, whereas the fracture strength is increased from 0.6 MPa to 3 MPa. 111We note that the rate of cross-link exchange in the quiescent vitrimers is relatively slow, as shown by the time-scale of the slow relaxation, as shown in Fig. S8; thus we expect that only if the rate of the cross-link exchange is driven by the applied deformation would it contribute to the stretchability.
Consecutive tensile loading-unloading cycles were also performed on the VMA-1’ sample with a waiting time of 10 min between each cycle. The results are shown for a maximum applied strain of 100% in Fig. S9 (SI) and 1000% in Fig. S10 (SI), respectively. No significant stress decrease is found for a repeated application of 100% strain, even though a small residual unrecovered strain () was observed right after the end of each cycle; this residual strain is similar to that previously reported ( 15%) during cyclic tensile testing of disulfide bond exchange-based PDMS vitrimers. [65]. In contrast, for large deformations (1000% strain), the required deformation stress is increasingly reduced for successive loading-unloading cycles, and a marked unrecovered strain () is observed, which takes more than 2 hours to recover.
The tensile response of all four semi-cured vitrimers are shown in Figure 5 (b). Beyond the yield stress, an increasing crosslink density corresponds to an increase in the stress required to deform the sample; an increase in the fracture stress from 0.6 MPa to 1.1 MPa; and a reduction from 8000% (VMA-1’) to 2750% (VMA-8’) of the elongation at break. Thus, while the tensile strength increases for increasing cross-link density, the elongation at break is reduced. Similarly, as shown in Figure 5(c), for the fully cured vitrimer samples, the tensile strength is considerably increased for increasing crosslink density, and even though the elongation at break is reduced compared with the semi-cured samples, it remains within the range of 1000% to 3000%, and the fracture stress for VMA-8 has a high value of 6.5 MPa. Although there have been a few recent reports of highly stretchable elastomers, [12, 10] our vitrimers uniquely combine very high stretchability with a high tensile strength, and importantly do not show the strain softening typically observed in other highly stretchable elastomers.[19, 20, 14]
II.5 Combination of toughness, strength and stretchability
A measure of mechanical toughness is obtained by integration of the stress-strain data in Figures 5(b–c), and the results are shown in Figure 5(d). The toughness generally increases upon full curing, with the exception of the VMA-1 sample for which a small decrease is observed, which is directly related to the reduction in the tensile extension achievable upon full curing. In Figure 5(e) the calculated toughness is also plotted versus the elongation at break both for the fully cured samples (highlighted by a yellow background) and for other recently reported examples of high performance vitrimers and reprocessable elastomers in the literature; the corresponding plot for the semi-cured samples is shown in Figure S11. Our vitrimers clearly show an excellent combination of stretchability and toughness, where the specific combination is tuneable by variation of the crosslink density. In fact, the toughness of our fully cured vitrimers is comparable to the best high performance elastomers in the literature, yet surpass these materials in stretchability, as shown in Figure 5(e). Moreover, our semi-cured vitrimers are comparable to other reported high-performance elastomers in the literature with regards to toughness, but can significantly outperform most other elastomers in terms of stretchability, see Figure S11.
To further illustrate the achieved materials properties, Figure 5(f) provides a radar chart for the fully cured samples with axes given by the elongation at break, the tensile strength at break, and the toughness, respectively. The combination of high toughness and elongation at break (particularly for VMA-2.5 and VMA-5) is noteworthy, as well as the trade-off between these properties and the tensile strength at break (as controlled by the crosslink density). The corresponding plot for the semicured samples is shown in Figure S12 (SI), demonstrating that semi-cured samples and particularly VMA-1’, combine an extremely high stretchability with excellent toughness and intermediate tensile strength at break; generally, the higher crosslink vitrimers gain in tensile strength at break, while showing a reduction in toughness and elongation at break.
II.6 Mechanical damping
The mechanical response, and particularly the mechanical damping, was investigated using tensile temperature ramp experiments performed at 1 Hz using a dynamic mechanical analyser (DMA). Representative results are shown for the VMA-1’ (dashed lines) and VMA-1 (solid lines) samples in Figure 5(g), where the storage (blue lines) and loss (green lines) are shown together with the loss tangent () (red lines). The main decay in the storage modulus, corresponding to a loss peak in , arises due to the structural () relaxation, as observed for the probe frequency of 1 Hz (corresponding to a relaxation time of 0.16 s; the smaller decay observed in at low temperatures, in turn corresponding to a weak peak in is due to a secondary relaxation, which persists within the glassy state.
Comparing these temperature-sweep DMA data with the SAOS frequency-sweep data of Fig. 4, it is clear that the -range where the ratio of loss to storage modulus, and thus , is the largest is found at temperatures just above the glass transition. Here, relaxations due to internal Rouse modes of the chain are superposed with a strong viscous damping due to the proximal glass transition, and the combination of these gives rise to a loss modulus significantly exceeding the storage modulus. In fact, this qualitative behaviour is seen between the glass transition and the entanglement/rubber plateau in polymeric materials in general; however, the separation of the loss modulus and the storage modulus is particularly large for polyacrylates. A plausible microscopic explanation for the effect is that dipole-dipole interactions between adjacent carbonyls (C=O C=O) in polyacrylates may lead to a large relaxation modulus associated with the glass transition, [60, 71] which significantly exceeds the entropic modulus of Rouse modes at the Kuhn segment length: this then leads to a greater viscous contribution of the glass transition within the Rouse relaxation region and so a higher , see Figure S14.
The -dependent loss tangent is plotted for all semicured (dashed lines) and fully cured (solid lines) vitrimer samples in Figure 5(h). A maximum loss factor () with a value between 2 and 3 is consistently observed. As the cross-link density increases upon full curing, the -relaxation slows down, as shown by the shift of the (Fig. 5g) and (Fig. 5h) loss peaks to higher temperatures, which is consistent with the increase in (see Table 1). Also, increasing the cross-link density, either by addition of cross-linker during synthesis or by curing, shifts the peak in the loss tangent, and both curing and nominal cross-linking concentration can thus be used to control the -position of the mechanical loss maximum. The temperature span of efficient mechanical loss ( 0.3) cover 35 5 °C, which is comparable to the ranges (°C) for commercial styrene-butadiene and natural rubbers ( 2.0). [72, 73] We note that the temperature range for efficient mechanical loss could be further increased by blending of semi-cured and fully cured vitrimers, thus producing materials with efficient loss factors across a range of 35 to 55 °C; such temperature ranges are comparable to those achieved for copolymer-based elastomers ( 1.7), [74] however, without the inherent compatibility issues that exist for blends of different rubbers ( 2.0). [75]
Since for damping applications, most (harmful) vibration sources are found in the frequency range of 0.01–100 Hz, [76, 35] the frequency-dependent shear mechanical response was also measured (see Experimental Section). The results are shown in Figure 5(i) for the fully cured vitrimers, demonstrating a maximum mechanical loss factor tan()2.5, in agreement with the single-frequency (1Hz) experiments (see panel (h)). Due to the restricted frequency window of the rheometer ( 3 decades), an empirical Havriliak-Negami expression [77] was used to describe the frequency-dependent loss tangent [78, 79], and thus provide a realistic estimate of the frequency-range for which the mechanical damping is efficient (tan()0.3, [80]) We conclude that our vitrimers demonstrate good mechanical damping across a frequency range of 0.001–100 Hz, corresponding to more than five orders of magnitude in frequency.
To put our results into perspective, the maximum value of the damping loss factor reported in literature is 1.0 by controlling the hydrogen bond properties in supramolecular networks, [81, 82, 83] 1.0 to 1.4 by use of microphase separation in long-fluorinated side-group polyacrylates or gels prepared by copolymerization of dual monomers with different solvent solubilities. [60, 84] It is 1.2 by utilising a mortise-and-tenon joint inspired rotaxane-containing mechanically interlocked supramolecular network, [59] around 1.0–1.5 for liquid crystal elastomers, [85, 86, 87] 1.5 by use of synergistic covalent and supramolecular interactions, [88] and close to 2.0 in polymer-fluid-based gels. [89] Thus, our vitrimers reach loss factor values at least as high as those for the best currently identified materials, and demonstrate effective loss factors across a wide ( 5 decades) and relevant frequency range for damping, spanning a wide -range around room temperature.
II.7 Room temperature self-healing
At elevated temperatures where the crosslink exchange kinetics are sufficiently fast, vitrimers are typically reprocessable. The dynamic nature of the network of covalent cross-links could also facilitate efficient self-healing if the exchange dynamics are well tuned to the operational temperature range. Efficient self-healing depends on both segmental mobility and bond exchange kinetics, in addition to interfacial properties such as wetting and adhesion. [90, 91, 92] Thus, for many investigated vitrimer systems, temperatures significantly above room temperature are required to achieve self-healing on realistic time-scales. [93, 94, 95] Previous work on a polymer network with crosslink exchange reactions based on boronic transesterification demonstrated room temperature self-healing with a healing time of 4 days; it was also shown that addition of water to the fracture surface accelerates self-healing as it promotes the formation of free boronic acid and diol groups at the interface, which can in turn lead to the formation of bonds across the interface. [32] These results, suggest that room temperature self-healing might be possible also in our PMA-based vitrimers, where the crosslink exchange is based on a similar mechanism.
To investigate the ability of our vitrimers to self-heal at room temperature (20 °C), the test procedure outlined in Figure 6(a) was followed. The tests were focused on the semicured VMA-1’ sample for which is lower than room temperature (20 °C), thus ensuring good segmental mobility at the fracture interface, The VMA-1’ sample was cut into two parts using a razor blade, after which the two parts were manually joined. To ensure that the two parts were fixed during healing, two glass slides and a small clip were used as an effective “splint” akin to the use of splints in the healing of bone fractures. Samples were prepared using both a 1 hour and a 24 hour healing time, respectively, and elongational stress-strain measurements were performed using an extensional speed of 50 mm/min (corresponding to a strain rate of 1.7 s-1), as shown in Figure 6(b). Data recorded for the original VMA-1’ sample (red) were compared with data recorded for the sample that had undergone healing for 1 hour (green) and 24 hours (blue), respectively. For the 1 hour sample, an elongation at fracture of 55 %, and a tensile strength at fracture of 74 %, of the corresponding values for the original sample were observed. After 24 hours, an elongation at fracture of 76 % and a tensile strength at fracture of 96 % were achieved.
The photos in Figure 6(c) show a representative VMA-1’ vitrimer sample before deformation at 0% strain (left), after 1 hour at a strain of 2500 (centre), and after 24 hours at a strain of 3000 (right). In the left and centre photos, the cut can still be observed, but after 24 hours of room temperature self-healing it is no longer easily distinguishable, as shown in the right photo. Thus, we conclude that our semi-cured vitrimers demonstrate good self-healing properties at room temperature, without the addition of water or any other ‘solvent’. We do note that due to the slower segmental dynamics (ca. 20 °C higher ) and slower (circa two orders of magnitude; see Figure S8) network rearrangments, the fully cured samples do not exhibit noticeable room temperature self-healing on a time-scale of 24 hours. However, fully cured vitrimers have the potential to self-heal at temperatures above room temperature.
III Conclusion
We here synthesize and characterise a new vitrimer system based on poly(methyl acrylate) with dynamic crosslinks utilising the associative exchange reactions of dioxaboralane metathesis. We demonstrate that these vitrimers combine ultra-stretchability and mechanical toughness with high mechanical dissipation and room temperature self-healing. Vitrimers with nominal cross-link densities between 2 and 8 mol% were prepared according to two different preparation routes: the first route involves a single curing stage (semi-cured) and the second an additional curing stage (fully cured); the two series of vitrimers show complementary properties with excellent mechanical and thermal stability over a wide temperature range (250 °C), and they can each be effectively re-processed using thermal or chemical processing conditions. The produced set of vitrimers demonstrate an excellent combination of ultra-stretchability (up to 8000 for semi-cured and 3000 for fully cured vitrimers) and mechanical toughness (20–55 MJ/m3), where the balance between the two is tuneable by variation of the crosslink density and the processing conditions. This combination of stretchability and toughness significantly outperforms most other elastomers in the literature, and the semi-cured vitrimers also demonstrate good self-healing at room temperature. Moreover, the vitrimers show excellent mechanical damping properties with a maximum mechanical loss factor value of 2.0–3.0, which corresponds to the values for the best currently identified damping materials, and shows a significant loss factor 0.3 across a wide (and for application relevant) frequency range of 5 decades (0.001–100 Hz), or correspondingly across a wide -range (35 5 °C for 1 Hz) around room temperature. Both the processing conditions and the variation of the nominal cross-link density can be used to tune the -range of efficient mechanical loss. Table S3 (SI) provides a comparison between the properties of our vitrimers and those of high performance elastomers in literature.
Our vitrimers are promising candidates in polymer-based applications where stretchability combined with mechanical toughness and/or a high mechanical dissipation is required. Examples include applications in soft robotics or biomimetic synthetic materials, such as artificial skin, muscles, or cartilage, where a combination of high stretching and self-healing ability, together with a high mechanical strength and dissipation are required. The vitrimers are also suited to applications such as protective devices (helmets or clothing) or in vibration dampers or shock absorbers. Future work will focus on further developing the detailed understanding of the links between the primary polymer chain length and dynamic cross-link distributions and the resulting vitrimer physical properties, and to expand these investigations to the wider family of polyacrylate-based vitrimers.
IV Experimental Section
IV.1 Materials
Methyl acrylate (99%, contains 100 ppm monomethyl ether hydroquinone as inhibitor), dichloromethane (DCM), 1,4-dioxane (anhydrous, 99.8%), deuterated dimethyl sulfoxide (DMSO-, 99.9 atom % D) and deuterium oxide (D2O, 99.9 atom % D) were purchased from Sigma Aldrich. 3-Allyloxy-1,2-Propanediol (99%) was purchased from Santa Cruz Biotechnology. Benzene-1 4-diboronic acid (98%) was purchased from Apollo Scientific Ltd. Magnesium sulfate (extra pure, dried), was purchased from Fisher Chemical. Azobisisobutyronitrile (AIBN) was purchased from VWR International Ltd. All the above reagents were used as received without further purification. Poly(methyl acrylate) (PMA-N315) was purchased from Sigma Aldrich, and this sample was supplied as a toluene solution. The solvent was first removed using a rotary evaporator, followed by three dissolution and reprecipitation cycles using THF and methanol to remove low molecular weight components to reach lower dispersity, and finally dried in a vacuum oven at 80 °C before use.
IV.2 Synthesis of diallyl dioxaborolane dynamic crosslinker
Benzene-1,4-diboronic acid (5.074 g, 30 mmol, 1 equiv.), 3-allyloxy-1,2-diol (7.50 ml, 60 mmol, 2 equiv.), dichloromethane (50 ml) and magnesium sulphate (10 g), used to absorb the water generated during the reaction, were added together in a 250 ml round-bottomed flask and stirred at room temperature for 3 hours. The solid magnesium sulphate was filtered off at the end of the reaction and a white solid product was obtained from the solution after removing the solvent using a rotary evaporator. 1H NMR (500 MHz, CDCl3): (ppm) = 7.83 (s, 4H, -Ph-), 5.88 (m, 1H, –CH=), 5.28 (dd, J = 17.3, 1.4 Hz, 1H, =CH2), 5.19 (dd, J = 10.4, 1.4 Hz, 1H, =CH2), 4.73 (m, 1H, –CH2–C(O–)H–CH2–), 4.43 (dd, J = 9.1, 8.2 Hz, 1H, –C(O–)H–CH2–O–), 4.18 (dd, J = 9.1, 6.6 Hz, 1H, –C(O–)H–CH2–O–), 4.06 (dq, J = 5.7, 1.6 Hz, 2H, –CH2–CH=), 3.64 (dd, J = 10.2, 5.1 Hz, 1H, –B–O–CH2–CH–), 3.57 (dd, J = 10.2, 5.1 Hz, 1H, –B–O–CH2–CH–). 13C NMR (125 MHz, CDCl3): (ppm) = 134.47 (–CH=CH2), 134.22 (–O–B(O–)C–CH–), 130.86 (–O–B(O–)C–), 117.60 (=CH2), 76.34 (-CH(CH2–)O–B–, 72.71 (=CH–CH2–), 72.03 (–B–O–CH2–), 68.62 (=CH–CH2–O–CH2–).
IV.3 Synthesis and preparation of PMA vitrimers
As an example, VMA-1 was prepared in a 100 ml Schlenk flask by adding previously synthesised diallyl dioxaborolane dynamic crosslinker 0.716 g (2 mmol), methyl acrylate 17.8 ml (198 mmol), initiator AIBN 0.164 g (1 mmol) and anhydrous dioxane 30 ml. Three freeze-pump-thaw cycles were performed before the reaction, and then the reaction was stirred at 70 °C under a nitrogen atmosphere until gelation. The reaction products were removed and left in a fume hood for two days to evaporate most of the solvents, and were then dried to constant weight in a vacuum oven at 80 °C. The dried product was placed in a mould and hot pressed at 120 °C using a pressure of 5 MPa for 15 minutes, then removed and naturally cooled to obtain the semi-cured VMA-1’ sample. To obtain the fully cured VMA-1 sample, the VMA-1’ sample was placed in a vacuum oven at 220 °C for 2 hours for secondary curing. Before testing, the fully cured vitrimers were hot pressed, as previously described, to prepare suitable sample geometries for the experiments.
IV.4 Nuclear magnetic resonance (NMR) spectroscopy
1H NMR spectra were recorded on a Bruker AV4 NEO 11.75 T 500 MHz spectrometer fitted with a 5 mm Bruker C/H cryoprobe. For the 1H NMR experiments, the spectral width was 9090 Hz, the relaxation delay time was 2 s, and the number of scans was 16. Around 0.5 mg VMA and 0.7 ml DMSO- were first added to an NMR tube, followed by the addition of 5 l of D2O to ensure hydrolysis of the boronic ester crosslinker. The samples were left for 48h after sample preparation and then measured to ensure that the hydrolysis was complete. Chemical shifts (in ppm) are referenced to the DMSO- solvent peak with a chemical shift of 2.50 ppm.
IV.5 Gel permeation chromatography (GPC)
The number average molecular weight (), weight average molecular weight (), and polymer dispersity () were measured using an Agilent 1260 Infinity system GPC instrument fitted with two 5 m Mixed-C columns plus a guard column, a refractive index detector, and a UV/Vis detector operating at 309 nm. THF (mixed with a small amount of water, vol/vol = 200:1) was used as the eluent (1 ml/min), and polystyrene standards ( ranging from 370 to 2,520,000 g/mol) were used for calibration. A polymer solution at a concentration of approximately 3 mg/ml was used for the measurements. The samples were left for 48h after sample preparation and then measured to ensure that hydrolysis was complete.
IV.6 Differential scanning calorimetry (DSC)
DSC measurements were performed using a TA Instruments Q2000 instrument equipped using a refrigerated cooling system. Samples with a typical weight of 5-10 mg were placed in hermetically sealed aluminium pans (Tzero pans from TA instruments). Experiments were performed between -70 °C and 100 °C and three cycles were run using a heating and cooling scan rate of 10 °C/min under a nitrogen flow of 50 ml/min.
IV.7 Thermogravimetric analysis (TGA)
TGA measurements were carried out using a TA Instruments Q50 TGA instrument under a nitrogen atmosphere from room temperature to 800 °C at a constant rate of 10 °C/min with a sample mass of 10 mg.
IV.8 Tensile Test
Tensile tests were performed room temperature (20 °C) using an Instron 5564 universal testing machine with a 2 kN capacity load cell. The tests were carried out on rectangular specimens (50 (length) 8 (width) 0.8 (thickness) mm) for varying extension speeds ranging from 10 to 200 mm/min.
IV.9 Dynamic mechanical analysis (DMA)
DMA experiments were performed using a TA Instruments Q850 DMA on vitrimer samples with sizes of around 10 mm (length) 4 mm (width) 0.5 mm (thickness). Temperature ramps from -50 to 100 °C were performed using a heating rate of 3 °C/min at a frequency of 1 Hz, using strain amplitudes within the linear viscoelastic region; a preload force of 1 N and a force track of 125% were used.
IV.10 Small-amplitude oscillatory shear (SAOS) rheology
SAOS was performed over a frequency range of 0.06-126 rad/s (0.01–20 Hz) using a Rheometrics ARES strain-controlled rheometer. A 5 mm diameter parallel plate geometry and a sample thickness of about 1.5 mm was used, and while cooling down, the sample thickness was gradually reduced (due to the increase in sample density) to maintain a constant sample radius. The temperature was controlled ( K) with a liquid nitrogen cooling system controlling a forced convection oven. Strain sweeps were performed prior to the frequency sweeps to confirm that for the strain amplitudes used, the measurements were performed in the linear viscoelastic region (LVR). The strain amplitude was decreased with increasing temperature to ensure that the measured torque remained within the relevant transducer range and that the measurements were performed within the LVR.
IV.11 Time temperature superposition of SAOS rheology data
Rheology data measured at different temperatures can often be combined into a mastercurve using so-called Time Temperature Superposition (TTS). However, TTS is only applicable when a single relaxation process controls the data, i.e. the data are rheologically simple. To investigate the relevance of TTS analysis, we plot the phase angle vs the complex modulus in a so-called van Gurp Palmen (vGP) plot [96], as shown for the VMA-8 sample in Fig. S15 (a). This representation of the data removes the explicit time-dependence and thus demonstrates whether TTS based on a simple horizontal shift parameter is a reasonable approximation. As shown in Fig. S5, the first observation is that for high moduli (low temperatures) a horizontal shifting works reasonably well, even though it is clear that generally a vertical shift parameter is also required. The vertical shift parameter for polymers is typically expressed as , where is the mass density. For our vitrimers, the detailed temperature variation of the density is unknown. Thus, we set as an approximation for the vertical shift parameter. To investigate the accuracy of this approach, we have plotted a modulus-renormalised vGP plot in Fig. S15 (b). The results confirm that a simple -renormalization provides a good approximation of the vertical shift.
Using the TA Instruments Orchestrator software, we first created mastercurves for the loss tangent () data, by determining the horizontal shift parameters required to overlap data for neighbouring temperatures; any vertical shifts required between data sets are removed for the loss tangent, thus facilitating an unbiased determination. Mastercurves for the shear storage and loss moduli, and , were subsequently created by also applying the vertical shift parameters , as described above. The resulting master curves for the shear storage and loss moduli, and are shown in Figure 4 for the four fully cured vitrimer samples, and in Fig. S7 (SI) for the semicured VMA-1’ sample. At low- the structural relaxation is the dominating mechanism, whereas at high-, a mechanism related to cross-link exchange is instead controlling the behaviour. Thus, we plot the mastercurves corresponding to these low and high temperature regimes separately in Fig. 4 and Fig. S7 (SI), respectively.
IV.12 Estimating the vitrimer mass density
The vitrimer samples were cut into well defined cuboids and weighed at room temperature (20 °C), which gave a mass density of 1.2 g/cm3. Given that the thickness of the vitrimer samples during SAOS experiments at 160 °C becomes ca 1.5 times that at 20 °C, and assuming that the sample cross-sectional areas do not change, the vitrimer densities at 160 °C were estimated to 0.8 g/cm3.
IV.13 Supporting Information
Supplementary Information to this article is available from.
IV.14 Acknowledgements
This work was supported by the EPSRC EP/T011726/1 and the Leverhulme Trust [RPG-2019-169]. Jiaxin Zhao thanks the China Scholarship Council and University of Leeds for awarding him a joint scholarship (202106220064).
IV.15 Conflict of Interest
The authors declare no conflict of interest.
IV.16 Data Availability Statement
The data in this paper are available in the Leeds Data Repository at https://doi.org/10.5518/1650.

### v2

@@
 Vitrimers are a relatively new class of polymer materials with unique properties offered by cross-links that can undergo associative exchange dynamics.
-We here present a new class of vitrimers based on poly(methyl acrylate) with cross-links utilising dioxaboralane metathesis.
+We here present a new class of vitrimers based on poly(methyl acrylate) with cross-links utilising dioxaborolane metathesis.
 These vitrimers demonstrate a combination of ultra-stretchability (up to 80 times their own length), mechanical toughness ( 40 MJ/m3), and thermal stability up to 250 °C; moreover, the vitrimers demonstrate excellent mechanical damping characterised by a loss factor () with a maximum of 2–3 and an effective value 0.3 across five decades in frequency (0.001–100 Hz), or correspondingly across a -range of 35 °C near room temperature (for a probe frequency of 1 Hz).
-The vitrimers can be successfully re-processed using both a thermo-mechanical and a chemical processing route, and can self-heal at room temperature, making them suitable for sustainable applications.
+The vitrimers can be successfully re-processed using both a thermo-mechanical and a chemical processing route, and can for low crosslink density self-heal at room temperature, making them suitable for sustainable applications.
 The material properties are directly tuneable by variation of both the amount of cross-linker and by the degree of curing.
@@
 A variety of approaches have been attempted often focused on combining an elastic stress-bearing network with a network of sacrificial bonds that can dissipate the applied mechanical energy by bond fracture.
-Reported polymer systems include: block-copolymer-based thermoplastic elastomers, [7] interpenetrating double polymer networks, [8, 9] elastomers based on supramolecular interactions, [10, 11] or polymeric networks based on dynamic covalent crosslinks, so called covalent adaptable networks (CANs) [12, 13, 14, 15, 16].
+Reported polymer systems include: block-copolymer-based thermoplastic elastomers, [7] interpenetrating double polymer networks, [8, 9] elastomers based on supramolecular interactions, [10, 11, 12] or polymeric networks based on dynamic covalent crosslinks, so called covalent adaptable networks (CANs) [13, 14, 15, 16, 17].
 However, elastomers that show very high stretchability ( 1000–10000%), typically also exhibit significant strain-softening upon yielding, leading to a rapidly decreasing mechanical strength upon deformation.
@@
 Here, we present a new ultra-stretchable vitrimer system based on poly(methyl acrylate) (PMA), where the PMA-based vitrimers are prepared by free radical polymerization (FRP) of the methyl acrylate monomer and a diallyl dioxaborolane dynamic crosslinker.
-The associative crosslink exhange reaction, based on dioxaboralane metathesis, shows excellent thermal stability and a relatively low activation energy (16 kJ/mol), which allows rapid bond exchange without added solvent, catalyst or free diols.
+The associative crosslink exchange reaction, based on dioxaborolane metathesis, shows excellent thermal stability and a relatively low activation energy (16 kJ/mol), which allows rapid bond exchange without added solvent, catalyst or free diols.
 [32] Long chain-length PMAs are viscous fluids at room temperature, characterised by glass-transition temperatures 15 °C, and are often used in coatings and adhesives, as liquid crystal elastomer backbone components, or as components in block co-polymers to improve material flexibility.
@@
 The successful polymerization of methyl acrylate is demonstrated by the disappearance of signals due to the monomer double bond (peaks a-c in Figure 1(c)).
-The NMR spectra demonstrate a 50-60% incorporation of the crosslinker after the initial FRP step (peaks labelled 1-3 in Figure 1(c)).
+The NMR spectra demonstrate a 50%–60% incorporation of the crosslinker after the initial FRP step (peaks labelled 1-3 in Figure 1(c)).
 The lower (than 100%) incorporation is due to the lower reactivity of the double bond in the allyl compared to the acrylate moiety.
@@
 In contrast, our vitrimers can be effectively re-processed using either a thermo-mechanical or a chemical route.
-For thermal reprocessing, the vitrimer samples were cut into pieces and placed into a hot-press mould, as shown in Figure 3(a)), where they were kept for 15 minutes at °C to facilitate dynamic cross-link exchange, and thus reprocessability.
+For thermal reprocessing, the vitrimer samples were cut into pieces and placed into a hot-press mould, as shown in Figure 3(a), where they were kept for 15 minutes at °C to facilitate dynamic cross-link exchange, and thus reprocessability.
 In contrast, the solvent-based reprocessing route is based on boronic ester hydrolysis, where the solvent needs to contain a sufficient amount of water for the hydrolysis (we used acetone mixed with a small amount of water; vol/vol = 200:1).
@@
 The plateau modulus can generally be attributed both to the covalent dynamic crosslinks and to topological crosslinks due to chain entanglements.
-In Figure S4, we compare the crosslink density , as determined from the plateau modulus (), with the nominal crosslink density determined from the chemical composition, showing that and demonstrating an approximately linear relation, even though a significant difference is observed between the magnitude of the nominal and actual crosslink density.
-The average molecular weight between crosslinks () can be estimated from , [52] where is the mass density 0.8 g/cm3 at 160 °C for VMA (see the detailed discussion in the Experimental Section).
-The estimated for VMA- (1, 2.5, 5 and 8) are: 80 kDa, 14 kDa, 6 kDa, and 4 kDa, respectively.
-For comparison, the entanglement molecular weight () for linear poly(methyl acrylate) (PMA) is 11 kDa, [44], so that is larger than at low crosslink density, and smaller than at high crosslink density.
+The average effective molecular weight between crosslinks () can be estimated from , [53] where is the mass density, which we estimate using the literature value g/cm3 for linear poly(methyl acrylate) at the relevant reference temperature of 160 °C [54].
+The resulting for VMA- (1, 2.5, 5 and 8) are: 109, 19, 8, and 5 kg/mol, respectively.
+For comparison, the entanglement molecular weight () for linear poly(methyl acrylate) (PMA) is 11 kg/mol, [45], so that is larger than at low crosslink density, and smaller than at high crosslink density.
 For context, it is also important to compare both and with the length of the highly polydisperse primary chains given in Table S1.
@@
 In all cases, due to the high polydispersity, we thus expect a significant number of chains that have only a small number of dynamic crosslinks, and/or which are shorter than the entanglement molecular weight.
-Both with regards to the significant difference between and , and the small influence of entanglements in determining the ‘rubbery’ plateau, we note that: (i) since the reactivity of the crosslinker is much lower than that of the acrylate monomer, we expect a degree of heterogeneity in the cross-link distribution, and the formation of “effective cross-links” (each containing several nominal cross-links) is not surprising; (ii) the calculation of the nominal cross-link density does not take account of the possibility of cross-link formation within the same chain; such ‘loops’ are not generally stress-bearing and will thus not contribute to the plateau modulus; (iii) it is highly likely, especially at low crosslink density, that many of the crosslinks lie within branched structures that are either “free” chains wholly unconnected to the network, or “side branches” connected to the network at one point only: such structures can relax their internal configuration, and so relax their contribution to the stress, without breaking of crosslinks and thus do not contribute to the plateau; (iv) finally, relaxation of such free chains or side branches will release entanglements through the well known process of dynamic dilution [51], thus progressively increasing the effective entanglement molecular weight and reducing the influence of entanglements on the plateau.
+Moreover, in Figure S4 we compare the estimated effective crosslink densities , as determined from the plateau moduli, with the corresponding nominal crosslink densities determined from the chemical composition and the mass density for the fully cured vitrimers.
+We determine using both the classical affine and phantom network models [53], respectively.
+Thus, the molar volumetric number density of stress-bearing chain strands was determined using , where for the affine network model, for the phantom network model, and is the cross-link functionality.
+We here use as an approximation that is consistent with the dynamic cross-linker chemistry [55], and for either model the cross-link density can be determined as .
+We find that for the affine network model the ratio lies within the range 0.04–0.12, and for the phantom network model within the range 0.08–0.25, for the investigated vitrimer samples.
+Thus, a significant difference between the estimated and the nominal crosslink density is consistently observed.
+Both with regards to the difference between and , and the small influence of entanglements in determining the ‘rubbery’ plateau, we note that: (i) since the reactivity of the crosslinker is much lower than that of the acrylate monomer, we expect a degree of heterogeneity in the cross-link distribution, and the formation of “effective cross-links” (each containing several nominal cross-links) is not surprising; (ii) the calculation of the nominal cross-link density does not take account of the possibility of cross-link formation within the same chain; such ‘loops’ are not generally stress-bearing and will thus not contribute to the plateau modulus; (iii) it is highly likely, especially at low crosslink density, that many of the crosslinks lie within branched structures that are either “free” chains wholly unconnected to the network, or “side branches” connected to the network at one point only: such structures can relax their internal configuration, and so relax their contribution to the stress, without breaking of crosslinks and thus do not contribute to the plateau; (iv) finally, relaxation of such free chains or side branches will release entanglements through the well known process of dynamic dilution [52], thus progressively increasing the effective entanglement molecular weight and reducing the influence of entanglements on the plateau.
 To further characterise the rheological response, we define two characteristic relaxation times from the crossover frequencies between the storage and loss moduli.
@@
 These values are slightly higher than those previously reported for vitrimers based on the same dynamic crosslinker, where examples include: polysiloxane vitrimers (20–57 kJ/mol), [56] vitrimers prepared by a thiol-ene “click” reaction (50 kJ/mol), [57] polyolefin vitrimers (8–47 kJ/mol), [58] and poly(styrene-b-isoprene-b-styrene) vitrimers (44 kJ/mol).
-[56] As shown in Figure 4(f) the activation energy increases monotonically with increasing nominal crosslink density, even though the increase for crosslink densities above 2.5 mol % is very small; this behaviour is consistent with reports for vitrimers based on both dioxaboralane and other exchange mechanisms.
+[59] As shown in Figure 4(f) the activation energy increases monotonically with increasing nominal crosslink density, even though the increase for crosslink densities above 2.5 mol% is very small; this behaviour is consistent with reports for vitrimers based on both dioxaborolane and other exchange mechanisms.
 [30, 55]
@@
 Thus, in the inset to Figure 5(a) we also present the estimated true stress () by assuming a constant volume during stretching.
-For the three lowest extension speeds, the true stress at break is effectively constant at 80 MPa, whereas at the slowest deformation, the true stress reaches only 50 MPa – hence, deformation at a faster rate results in the vitrimer reaching its ‘true’ strength at break, and thus fracture, sooner.
+For the three higher extension speeds, the true stress at break is effectively constant at 80 MPa, whereas at the slowest deformation, the true stress reaches only 50 MPa – hence, deformation at a faster rate results in the vitrimer reaching its ‘true’ strength at break, and thus fracture, sooner.
 Macroscopic failure of materials commonly starts with micro-cracks, where crack propagation leads to the final failure.
@@
 Thus, particularly for the slowest deformation rate, there is enough time for structural relaxation to take place during the deformation.
-It is also possible that the rate of the relaxation and/or the bond exchange mechanism is significantly sped up, particularly where the stress is focused near the tip of any micro-cracks, resulting in further contributions to these effects.
 The overall result is a negligible strain-hardening for the lowest applied strain rate and a remarkably stretchable material, for which a deformation at break of 8000% is observed; the corresponding deformation at break is reduced to 2500% for the faster rate of 3.024 s-1, whereas the fracture strength is increased from 0.6 MPa to 3 MPa.
-111We note that the rate of cross-link exchange in the quiescent vitrimers is relatively slow, as shown by the time-scale of the slow relaxation, as shown in Fig.
-S8; thus we expect that only if the rate of the cross-link exchange is driven by the applied deformation would it contribute to the stretchability.
+We note that it is possible that the rate of the relaxation and/or the bond exchange mechanism becomes significantly sped up during the deformation, particularly where the stress is focused near the tip of any micro-cracks, resulting in further contributions to the observed behaviour.
+The rate of cross-link exchange in the quiescent vitrimers is relatively slow, as shown by the time-scale of the network relaxation (see Fig.
+S8).
+Thus, we expect that only if this rate is significantly increased due to the applied deformation, will the dynamic nature of the crosslinks influence the stretchability.
+From our experiments alone, we cannot determine the role (if any) of the dynamic cross-links in the tensile deformation behaviour.
+However, previous studies on PMA-based elastomers with permanent crosslinks [72, 73] have also demonstrated highly stretchable behaviour, with a maximum stretchability of 1200% [72] observed for the lowest investigated crosslink density ( mol/cm3) and strain rate (0.01 s-1).
+This does not exclude a role of the dynamic cross-links, but it does illustrate the key role played by the near room temperature location of the glass transition for PMA in providing a highly stretchable material.
 Consecutive tensile loading-unloading cycles were also performed on the VMA-1’ sample with a waiting time of 10 min between each cycle.
@@
 5g) and (Fig.
-5h) loss peaks to higher temperatures, which is consistent with the increase in (see Table 1).
+5h) loss peaks to higher temperatures, which is consistent with the increase in (see Table S2 in the SI).
 Also, increasing the cross-link density, either by addition of cross-linker during synthesis or by curing, shifts the peak in the loss tangent, and both curing and nominal cross-linking concentration can thus be used to control the -position of the mechanical loss maximum.
@@
 [92] Thus, our vitrimers reach loss factor values at least as high as those for the best currently identified materials, and demonstrate effective loss factors across a wide ( 5 decades) and relevant frequency range for damping, spanning a wide -range around room temperature.
-II.7 Room temperature self-healing
+II.7 Self-healing
 At elevated temperatures where the crosslink exchange kinetics are sufficiently fast, vitrimers are typically reprocessable.
@@
 We do note that due to the slower segmental dynamics (ca.
-20 °C higher ) and slower (circa two orders of magnitude; see Figure S8) network rearrangments, the fully cured samples do not exhibit noticeable room temperature self-healing on a time-scale of 24 hours.
+20 °C higher ) and slower (circa two orders of magnitude; see Figure S8) network rearrangements, the fully cured samples do not exhibit noticeable room temperature self-healing on a time-scale of 24 hours.
 However, fully cured vitrimers have the potential to self-heal at temperatures above room temperature.
 III Conclusion
-We here synthesize and characterise a new vitrimer system based on poly(methyl acrylate) with dynamic crosslinks utilising the associative exchange reactions of dioxaboralane metathesis.
-We demonstrate that these vitrimers combine ultra-stretchability and mechanical toughness with high mechanical dissipation and room temperature self-healing.
-Vitrimers with nominal cross-link densities between 2 and 8 mol% were prepared according to two different preparation routes: the first route involves a single curing stage (semi-cured) and the second an additional curing stage (fully cured); the two series of vitrimers show complementary properties with excellent mechanical and thermal stability over a wide temperature range (250 °C), and they can each be effectively re-processed using thermal or chemical processing conditions.
-The produced set of vitrimers demonstrate an excellent combination of ultra-stretchability (up to 8000 for semi-cured and 3000 for fully cured vitrimers) and mechanical toughness (20–55 MJ/m3), where the balance between the two is tuneable by variation of the crosslink density and the processing conditions.
-This combination of stretchability and toughness significantly outperforms most other elastomers in the literature, and the semi-cured vitrimers also demonstrate good self-healing at room temperature.
+We have synthesized and characterised a new vitrimer system based on poly(methyl acrylate) with dynamic crosslinks utilising the associative exchange reactions of dioxaborolane metathesis.
+We demonstrate that these vitrimers combine ultra-stretchability and mechanical toughness with high mechanical dissipation and self-healing.
+Vitrimers with nominal cross-link densities between 2 and 8 mol% were prepared according to two different preparation routes: the first route involved a single curing stage (semi-cured) and the second an additional curing stage (fully cured); the two series of vitrimers showed complementary properties with excellent mechanical and thermal stability over a wide temperature range (250 °C), and they could each be effectively re-processed using thermal or chemical processing conditions.
+The produced set of vitrimers demonstrated an excellent combination of ultra-stretchability (up to 8000 for semi-cured and 3000 for fully cured vitrimers) and mechanical toughness (20–55 MJ/m3), where the balance between the two is tuneable by variation of the crosslink density and the processing conditions.
+This combination of stretchability and toughness significantly outperforms most other elastomers in the literature, and the lowest cross-link semi-cured vitrimers also demonstrate good self-healing at room temperature.
 Moreover, the vitrimers show excellent mechanical damping properties with a maximum mechanical loss factor value of 2.0–3.0, which corresponds to the values for the best currently identified damping materials, and shows a significant loss factor 0.3 across a wide (and for application relevant) frequency range of 5 decades (0.001–100 Hz), or correspondingly across a wide -range (35 5 °C for 1 Hz) around room temperature.
@@
 The vitrimers are also suited to applications such as protective devices (helmets or clothing) or in vibration dampers or shock absorbers.
-Future work will focus on further developing the detailed understanding of the links between the primary polymer chain length and dynamic cross-link distributions and the resulting vitrimer physical properties, and to expand these investigations to the wider family of polyacrylate-based vitrimers.
+Future work will focus on further developing the detailed understanding of the links between the primary polymer chain length, the detailed dynamics and the distribution of cross-link, and the resulting vitrimer physical properties, as well as to expand these investigations to the wider family of polyacrylate-based vitrimers.
 IV Experimental Section
@@
 Around 0.5 mg VMA and 0.7 ml DMSO- were first added to an NMR tube, followed by the addition of 5 l of D2O to ensure hydrolysis of the boronic ester crosslinker.
-The samples were left for 48h after sample preparation and then measured to ensure that the hydrolysis was complete.
+The samples were left for 48 hours after sample preparation and then measured to ensure that the hydrolysis was complete.
 Chemical shifts (in ppm) are referenced to the DMSO- solvent peak with a chemical shift of 2.50 ppm.
@@
 A polymer solution at a concentration of approximately 3 mg/ml was used for the measurements.
-The samples were left for 48h after sample preparation and then measured to ensure that hydrolysis was complete.
+The samples were left for 48 hours after sample preparation and then measured to ensure that hydrolysis was complete.
 IV.6 Differential scanning calorimetry (DSC)
@@
 S7 (SI), respectively.
-IV.12 Estimating the vitrimer mass density
-The vitrimer samples were cut into well defined cuboids and weighed at room temperature (20 °C), which gave a mass density of 1.2 g/cm3.
-Given that the thickness of the vitrimer samples during SAOS experiments at 160 °C becomes ca 1.5 times that at 20 °C, and assuming that the sample cross-sectional areas do not change, the vitrimer densities at 160 °C were estimated to 0.8 g/cm3.
-IV.13 Supporting Information
+IV.12 Supporting Information
 Supplementary Information to this article is available from.
-IV.14 Acknowledgements
+IV.13 Acknowledgements
 This work was supported by the EPSRC EP/T011726/1 and the Leverhulme Trust [RPG-2019-169].
 Jiaxin Zhao thanks the China Scholarship Council and University of Leeds for awarding him a joint scholarship (202106220064).
-IV.15 Conflict of Interest
+IV.14 Conflict of Interest
 The authors declare no conflict of interest.
-IV.16 Data Availability Statement
+IV.15 Data Availability Statement
 The data in this paper are available in the Leeds Data Repository at https://doi.org/10.5518/1650.